OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

I don't quite understand what its asking for ... Do you mean like: a. 45 = (7-2)(7+2) b. 77 = (9-2)(9+2) c. 112 = (11-3)(11)

OpenStudy (anonymous):

i'm not sure